OpenStudy (anonymous):

Use rational exponents to write ** as a single radical expression (will input in this link).

OpenStudy (anonymous):



OpenStudy (saifoo.khan):

that's blank/.

OpenStudy (anonymous):

My computers dumb...its a tiny exponent 3, square root symbol over a large 6 times tiny exponent 2, square root symbol over a large 7

OpenStudy (anonymous):

AAAAHHHHH

OpenStudy (saifoo.khan):

can use use the equation editor?

OpenStudy (anonymous):

\[\sqrt[2]{7} \sqrt[3]{6} \]

OpenStudy (anonymous):

multiplied

OpenStudy (saifoo.khan):

4.8

OpenStudy (anonymous):

Use rational exponents to write ** as a single radical expression

OpenStudy (saifoo.khan):

\[\Large 7^\frac12 \times 6^\frac13\]

OpenStudy (saifoo.khan):

\[\Huge 42^\frac56\]

OpenStudy (saifoo.khan):

\[\huge \sqrt[6]{42^5}\]